# Classical harmonic analysis viewed through the prism of noncommutative geometry

### v3

@@ 4 Coupling between Banach K-homology with K-theory
While Lafforgue’s $\mathrm{K}\mathrm{K}^{\mathrm{ban}}$ provides a powerful bivariant framework and an abstract product formalism [Laf02], it does not supply a concrete Toeplitz-type formula for the pairing. The novelty here is precisely this explicit, calculable index formula in a non-Hilbert Banach context. The construction applies uniformly to classes $\mathscr{B}$ stable under isomorphisms and finite sums (e.g., the class of $\mathrm{L}^{p}$-spaces), and it will serve as the analytic backbone for the Banach-space Chern characters developed later in the paper.
@@
Let $\mathcal{A}$ be a unital algebra and let $\mathscr{B}$ be a class of Banach spaces stable under isomorphisms and finite sums and containing the zero space. Consider an invertible $u$ of the matrix algebra $\mathrm{M}_{n}(\mathcal{A})$ and an odd Banach Fredholm module $(X,\pi,F)$ over $\mathcal{A}$ with $\pi(1)=\mathrm{Id}_{X}$. We introduce the bounded operators $P_{n}\overset{\mathrm{def}}{=}\mathrm{Id}\otimes\frac{\mathrm{Id}+F}{2}\colon X^{\oplus n}\to X^{\oplus n}$ and $u_{n}\overset{\mathrm{def}}{=}(\mathrm{Id}\otimes\pi)(u)\colon X^{\oplus n}\to X^{\oplus n}$. Then the bounded operator
@@
Let $\mathcal{A}$ be a unital algebra and let $\mathscr{B}$ be a class of Banach spaces stable under isomorphisms and finite sums and containing the zero space. Consider a projection $e$ of the matrix algebra $\mathrm{M}_{n}(\mathcal{A})$ and an even Banach Fredholm module $\bigg(X_{+}\oplus X_{-},\pi,\begin{bmatrix}0&F_{-}\\ F_{+}&0\\ \end{bmatrix},\gamma\bigg)$ over $\mathcal{A}$ with $\pi(1)=\mathrm{Id}_{X}$. We introduce the bounded operator $e_{n}\overset{\mathrm{def}}{=}(\mathrm{Id}\otimes\pi)(e)\colon X^{\oplus n}\to X^{\oplus n}$. Then the bounded operator
@@ 5 Summability of Banach Fredholm modules
This concept was introduced by Pietsch in [Pie74]. If $X$ and $Y$ are Hilbert spaces and if $T\colon X\to Y$ is a compact operator then by [Kon01, p. 69] these numbers coincide with the singular numbers of the operator $T$. We refer to the books [Pi80], [Pie87], [Kon01] and to the survey papers [Kon01], [Ger25] and [Pie23] for more information. The history of this topic is equally described in [Pie07, Chapter 6]. Let us now describe a few examples.
@@
$a_{n}(T)\overset{\mathrm{def}}{=}\inf\{\left\|T-R\right\|:R\colon X\to Y,\,\operatorname{rank}R<n\},\quad n\geqslant 1.$
@@
Actually, the approximation numbers are the largest $s$-numbers by [Kon86, Lemma p. 69]. Finally, when $Y$ is reflexive, by [EdT86, Remark 7 p. 237] we have
(5.4) $a_{n}(T)=a_{n}(T^{*}),\quad n\geqslant 1.$
@@
Suppose that $0<p<\infty$ and $0<q\leqslant\infty$. Let $T\mapsto(s_{n}(T))_{n\geqslant 0}$ be an $s$-number sequence. For any Banach spaces $X$ and $Y$, following [Kon86, Definition 1.d.18 p. 72] we define the classes
(5.6) $S^{p,q}_{s}(X,Y)\overset{\mathrm{def}}{=}\big\{T\colon X\to Y:(s_{n}(T))\in\ell^{p,q}\big\}$
@@
Moreover, if $T\in S^{p,q}_{s}(X,Y)$, we let
$\left\|T\right\|_{S^{p,q}_{s}(X,Y)}\overset{\mathrm{def}}{=}\left\|(s_{n}(T))\right\|_{\ell^{p,q}}.$
By [Kon86, Lemma p. 72], $(S^{p,q}_{s},\left\|\cdot\right\|_{S^{p,q}_{s}})$ is a quasi-Banach ideal. For the cases $s_{n}=a_{n}$ or if $s_{n}=x_{n}$ we will use the notations $S^{p,q}_{\mathrm{app}}$ and $S^{p,q}_{\mathrm{weyl}}$. If $X=Y$, we use the notation $S^{p,q}_{s}(X)$. If, in addition $X$ is equal to a Hilbert space $H$, we recover the Schatten class $S^{q}(H)$ with the space $S^{q}(X)$.
If $0<p<q<\infty$, we have obviously the contractive inclusion $S^{p}_{s}(X)\subset S^{q}_{s}(X)$. Furthermore, if $\frac{1}{r}=\frac{1}{p}+\frac{1}{q}$, we have by [Kon86, Proposition 1.d.19 p. 73] the inclusion
@@ 6 The Chern character of a $q$-summable Banach Fredhlom module
$=P([P,u^{-1}])uP=\frac{1}{2}P\mathop{}\mathopen{}\mathrm{d}u^{-1}uP=\frac{1}{2}P\mathop{}\mathopen{}\mathrm{d}u^{-1}(uP-Pu+Pu)$
@@ 7 From Banach spectral triples to Banach Fredhlom modules
We refer to [ArK22, Section 5.7] for some properties of $\operatorname{\mathrm{Lip}}_{D}(\mathcal{A})$. The following definition is essentially [ArK22, Definition 5.10 p. 218]. Here $\rho(D)\overset{\mathrm{def}}{=}\mathbb{C}\backslash\sigma(D)$ is the resolvent set of the operator $D$.
@@
Consider a triple $(\mathcal{A},X,D)$ constituted of the following data: a reflexive3332. It may perhaps be possible to replace the reflexivity by an assumption of weak compactness, see [HvNVW18, p. 361]. Banach space $X$, a bisectorial operator $D$ on $X$ with dense domain $\operatorname{dom}D\subset X$, and a Banach algebra $\mathcal{A}$ equipped with a continuous homomorphism $\pi\colon\mathcal{A}\to\mathrm{B}(X)$. We say that $(\mathcal{A},X,D)$ is a compact Banach spectral triple if
@@
    the subset $\operatorname{\mathrm{Lip}}_{D}(A)$ is dense in the algebra $\mathcal{A}$.
@@
    The subset $\operatorname{\mathrm{Lip}}_{D}(A)$ is dense in the algebra $\mathcal{A}$.
@@
The operator $D$ is bisectorial. Consequently, by Proposition 2.10 the operator $D^{2}$ is sectorial. Moreover, since $D$ admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\theta}^{\mathrm{bi}})$ functional calculus, $D^{2}$ admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\theta})$ functional calculus by [HvNVW18]. Recall that the functional calculus of the operators $D$ and $D^{2}$ are compatible. So using the change of variables $t=s^{2}$ in the last equality, we obtain for any $x\in\operatorname{Ran}D$
@@
Let $(\mathcal{A},X,D)$ be a $q$-summable compact Banach spectral triple for some $q>0$. If $S^{q}_{\mathrm{app}}(X)$ has the strong convex compactness property then the proof of Proposition 7.7 shows that the Banach Fredholm module $(X,\pi,\operatorname{\mathrm{sgn}}D)$ is $q$-summable.
@@ 8 Examples of triples and Banach Fredholm modules
Suppose that $1<p<\infty$. Here $\mathbb{T}\overset{\mathrm{def}}{=}\{z\in\mathbb{C}:|z|=1\}$ is the one-dimensional torus. The functions defined on $\mathbb{T}$ can be identified with periodic functions on $\mathbb{R}$ with period $2\pi$. In particular, we have
$\mathrm{C}(\mathbb{T})=\{g\in\mathrm{C}(\mathbb{R}):f(\theta+2\pi)=f(\theta)\}$
Recall that by [EnN00, Proposition 4.10 p. 31] any multiplication operator $M_{f}\colon\mathrm{L}^{p}(\mathbb{T})\to\mathrm{L}^{p}(\mathbb{T})$, $g\mapsto fg$ by a measurable function $f\colon\mathbb{T}\to\mathbb{C}$ is bounded if and only if $f$ is essentially bounded. In this case, we have $\left\|M_{f}\right\|_{\mathrm{L}^{p}(\mathbb{T})\to\mathrm{L}^{p}(\mathbb{T})}=\mathrm{L}^{\infty}(\mathbb{T})$. Consequently, we can consider the isometric homomorphism $\pi\colon\mathrm{L}^{\infty}(\mathbb{T})\to\mathrm{B}(\mathrm{L}^{p}(\mathbb{T}))$, $f\mapsto M_{f}$. It is well known, essentially from [Kha13, Exercise 4.2.6 p. 194] and [GVF01, p. 390], that $(\mathrm{C}(\mathbb{T}),\mathrm{L}^{2}(\mathbb{T}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta})$ is a compact spectral triple, where the used homomorphism is the restriction of $\pi$ on the algebra $\mathrm{C}(\mathbb{T})$. First, we prove an $\mathrm{L}^{p}$-generalization of this classical fact. Here, we consider the closure of the closable4443. Let $(f_{n})_{n\geqslant 1}$ be a sequence of functions belonging to $\mathrm{C}^{\infty}(\mathbb{T})$ such that $f_{n}\to 0$ in $\mathrm{L}^{p}(\mathbb{T})$ and $f_{n}^{\prime}\to g$ in $\mathrm{L}^{p}(\mathbb{T})$ for some $g\in\mathrm{L}^{p}(\mathbb{T})$. For any $h\in\mathrm{C}^{\infty}(\mathbb{T})$, integration by parts gives $\int_{0}^{2\pi}f_{n}^{\prime}(\theta)h(\theta)\mathop{}\mathopen{}\mathrm{d}\theta=-\int_{0}^{2\pi}f_{n}(\theta)h^{\prime}(\theta)\mathop{}\mathopen{}\mathrm{d}\theta$. Passing to the limit (using Hölder’s inequality), we obtain $\int_{0}^{2\pi}g(\theta)h(\theta)\mathop{}\mathopen{}\mathrm{d}\theta=0$ for all $h\in\mathrm{C}^{\infty}(\mathbb{T})$. Hence $g=0$ in $\mathrm{L}^{p}(\mathbb{T})$. unbounded operator $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}\colon\mathrm{C}^{\infty}(\mathbb{T})\subset\mathrm{L}^{p}(\mathbb{T})\to\mathrm{L}^{p}(\mathbb{T})$ and we denote again $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ its closure whose the domain $\operatorname{dom}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ is the Sobolev space
@@
Note that this operator is an unbounded Fourier multiplier with symbol $(n)_{n\in\mathbb{Z}}$.
If $\alpha\geqslant 0$, we also need to consider the fractional Sobolev space
(8.1) $\mathrm{L}^{p}_{\alpha}(\mathbb{T})\overset{\mathrm{def}}{=}\operatorname{dom}\Delta^{\frac{\alpha}{2}}_{p},$
which is a subspace of the Banach space $\mathrm{L}^{p}(\mathbb{T})$. For any function $f\in\mathrm{L}^{p}_{\alpha}(\mathbb{T})$, we will use the notation
(8.2) $\left\|f\right\|_{\mathrm{L}^{p}_{\alpha}(\mathbb{T})}\overset{\mathrm{def}}{=}\big\|(\mathrm{Id}+\Delta_{p})^{\frac{\alpha}{2}}(f)\big\|_{\mathrm{L}^{p}(\mathbb{T})}.$
In particular, we have $\mathrm{L}^{p}_{1}(\mathbb{T})=\mathrm{W}^{1,p}(\mathbb{T})$. We refer to [Str83] for the treatment of the more general case of fractional Sobolev space on complete Riemannian manifolds.
@@
(8.3) $\left\|f\right\|_{\mathrm{L}^{q}(\mathbb{T}^{d})}\lesssim\left\|f\right\|_{\mathrm{L}^{p}_{\alpha}(\mathbb{T}^{d})}.$

### v4

@@ 4 Coupling between Banach K-homology with K-theory
While Lafforgue’s $\mathrm{K}\mathrm{K}^{\mathrm{ban}}$ provides a powerful bivariant framework and an abstract product formalism [Laf02], it does not supply a concrete Toeplitz-type formula for the pairing. The novelty here is precisely this explicit, calculable index formula in a non-Hilbert Banach context. The construction applies uniformly to classes $\mathscr{B}$ stable finite sums (e.g., the class of $\mathrm{L}^{p}$-spaces up to isomorphism), and it will serve as the analytic backbone for the Banach-space Chern characters developed later in the paper.
@@
Let $\mathcal{A}$ be a unital algebra and let $\mathscr{B}$ be a class of Banach spaces stable under finite sums and containing the zero space. Consider an invertible $u$ of the matrix algebra $\mathrm{M}_{n}(\mathcal{A})$ and an odd Banach Fredholm module $(X,\pi,F)$ over $\mathcal{A}$ with $\pi(1)=\mathrm{Id}_{X}$. We introduce the bounded operators $P_{n}\overset{\mathrm{def}}{=}\mathrm{Id}\otimes\frac{\mathrm{Id}+F}{2}\colon X^{\oplus n}\to X^{\oplus n}$ and $u_{n}\overset{\mathrm{def}}{=}(\mathrm{Id}\otimes\pi)(u)\colon X^{\oplus n}\to X^{\oplus n}$. Then the bounded operator
@@
Let $\mathcal{A}$ be a unital algebra and let $\mathscr{B}$ be a class of Banach spaces stable under finite sums and containing the zero space. Consider a projection $e$ of the matrix algebra $\mathrm{M}_{n}(\mathcal{A})$ and an even Banach Fredholm module $\bigg(X_{+}\oplus X_{-},\pi,\begin{bmatrix}0&F_{-}\\ F_{+}&0\\ \end{bmatrix},\gamma\bigg)$ over $\mathcal{A}$ with $\pi(1)=\mathrm{Id}_{X}$. We introduce the bounded operator $e_{n}\overset{\mathrm{def}}{=}(\mathrm{Id}\otimes\pi)(e)\colon X^{\oplus n}\to X^{\oplus n}$. Then the bounded operator
@@ 5 Summability of Banach Fredholm modules
This concept was introduced by Pietsch in [Pie74]. If $X$ and $Y$ are Hilbert spaces and if $T\colon X\to Y$ is a compact operator then by [Kon01, p. 69] these numbers coincide with the singular numbers of the operator $T$. We refer to the books [Pi80], [Pie87], [Kon01] and to the survey papers [Kon01], [Ger25] and [Pie23] for more information. The history of this topic is equally described in [Pie07, Chapter 6]. The $s$-numbers are continuous functions since $|s_{n}(S)-s_{n}(T)|\leqslant\left\|S-T\right\|$ for any bounded operators $S,T\colon X\to Y$ by [Pie74, Theorem 1 p. 203]. Finally, when $Y$ is reflexive we have by [EdT86, Remark 7 p. 237] the equality
(5.2) $a_{n}(T)=a_{n}(T^{*}),\quad n\geqslant 1.$
Let us now describe a few examples.
@@
(5.3) $a_{n}(T)\overset{\mathrm{def}}{=}\inf\{\left\|T-R\right\|:R\colon X\to Y,\,\operatorname{rank}R<n\},\quad n\geqslant 1.$
@@
Actually, the approximation numbers are the largest $s$-numbers by [Kon86, Lemma p. 69] or [Pi80, Theorem 11.2.3 p. 146].
We refer to [Pie07, Section 6.2] for more examples of $s$-number sequences.
@@
Suppose that $0<q<\infty$ and $0<r\leqslant\infty$. Let $T\mapsto(s_{n}(T))_{n\geqslant 0}$ be an $s$-number sequence. For any Banach spaces $X$ and $Y$, following [Kon86, Definition 1.d.18 p. 72] we define the classes
(5.7) $S^{q,r}_{s}(X,Y)\overset{\mathrm{def}}{=}\big\{T\colon X\to Y:(s_{n}(T))\in\ell^{q,r}\big\}$
@@
Moreover, if $T\in S^{q,r}_{s}(X,Y)$, we let
(5.9) $\left\|T\right\|_{S^{q,r}_{s}(X,Y)}\overset{\mathrm{def}}{=}\left\|(s_{n}(T))\right\|_{\ell^{q,r}}=\begin{cases}\displaystyle\bigg(\sum_{n=1}^{\infty}n^{\frac{r}{q}-1}s_{n}(T)^{r}\bigg)^{\frac{1}{r}}&\text{if }0<r<\infty\\ \displaystyle\sup_{n\geqslant 1}n^{\frac{1}{q}}s_{n}(T)&\text{if }r=\infty\\ \end{cases}.$
By [Kon86, Lemma p. 72], $(S^{q,r}_{s},\left\|\cdot\right\|_{S^{q,r}_{s}})$ is a quasi-Banach ideal. For the cases $s_{n}=a_{n}$ or if $s_{n}=x_{n}$ we will use the notations $S^{q,r}_{\mathrm{app}}$ and $S^{q,r}_{\mathrm{weyl}}$. If $X=Y$, we use the notation $S^{q,r}_{s}(X)$. If, in addition $X$ is equal to a Hilbert space $H$, we recover the Schatten class $S^{q}(H)$ with the space $S^{q}(X)$.
If $0<p<q<\infty$, we have obviously the contractive inclusion $S^{p}_{s}(X)\subset S^{q}_{s}(X)$ and $S^{q}_{s}(X)\subset S^{q,\infty}_{s}(X)$. Furthermore, if $\frac{1}{r}=\frac{1}{p}+\frac{1}{q}$, we have by [Kon86, Proposition 1.d.19 p. 73] the inclusion
@@
Finally, if the Banach space $X$ has the approximation property, then by [Pi80] we have $s_{n}(T)\to 0$ if and only if $T$ is compact.
@@ 6 The Chern character of a $q$-summable Banach Fredhlom module
$=P([P,u^{-1}])uP=\frac{1}{2}P\mathop{}\mathopen{}\mathrm{d}u^{-1}uP$
$=\frac{1}{2}P\mathop{}\mathopen{}\mathrm{d}u^{-1}(uP-Pu+Pu)$
@@
Consequently, this operator also belongs to the space $S^{N}_{\mathrm{app}}(X)$. So we use Corollary 6.5 with the operators $PuP\colon P(X)\to P(X)$ and $Pu^{-1}P\colon P(X)\to P(X)$ replacing $T$ and $R$. Using the equalities
$\operatorname{Tr}(\mathop{}\mathopen{}\mathrm{d}u^{-1}\mathop{}\mathopen{}\mathrm{d}u)^{N}=\operatorname{Tr}(\mathop{}\mathopen{}\mathrm{d}u\mathop{}\mathopen{}\mathrm{d}u^{-1})^{N}$
and
@@ 7 From Banach spectral triples to Banach Fredhlom modules
We refer to [ArK22, Section 5.7] for some properties of $\operatorname{\mathrm{Lip}}_{D}(\mathcal{A})$. If $\mathcal{B}$ is a subalgebra of $\mathcal{A}$, observe that we have $\operatorname{\mathrm{Lip}}_{D}(\mathcal{B})=\operatorname{\mathrm{Lip}}_{D}(\mathcal{A})\cap\mathcal{B}$.
The following definition is essentially [ArK22, Definition 5.10 p. 218]. Here $\rho(D)\overset{\mathrm{def}}{=}\mathbb{C}\backslash\sigma(D)$ is the resolvent set of the operator $D$.
@@
Consider a triple $(\mathcal{A},X,D)$ constituted of the following data: a reflexive4443. It may perhaps be possible to replace the reflexivity by an assumption of weak compactness, see [HvNVW18, p. 361]. Banach space $X$, a bisectorial operator $D$ on $X$ with dense domain $\operatorname{dom}D\subset X$, and a Banach algebra $\mathcal{A}$ equipped with a continuous homomorphism $\pi\colon\mathcal{A}\to\mathrm{B}(X)$. We say that $(\mathcal{A},X,D)$ is a compact Banach spectral triple if
@@
    the subset $\operatorname{\mathrm{Lip}}_{D}(\mathcal{A})$ is dense in the algebra $\mathcal{A}$.
@@
    The subset $\operatorname{\mathrm{Lip}}_{D}(\mathcal{A})$ is dense in the algebra $\mathcal{A}$.
@@
In this case, the spectral dimension of the locally compact Banach spectral triple is defined by
(7.4) $\dim(\mathcal{A},X,D)\overset{\mathrm{def}}{=}\inf\big\{q>0:\text{for any $a\in\operatorname{\mathrm{Lip}}_{D}(\mathcal{A})$ we have }\pi(a)(D+\mathrm{i})^{-1}\in S^{q,\infty}_{\mathrm{app}}(X)\big\}.$
In the degenerated situation, we need to replace $(D+\mathrm{i})^{-1}$ by $|D|^{-1}$.
@@
The operator $D$ is bisectorial and admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\theta}^{\mathrm{bi}})$ functional calculus. Consequently, by Proposition 2.10 the operator $D^{2}$ is sectorial and admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\theta})$ functional calculus. Recall that the functional calculus of the operators $D$ and $D^{2}$ are compatible. So using the change of variables $t=s^{2}$ in the last equality, we obtain for any $x\in\operatorname{Ran}D$
@@
Let $(\mathcal{A},X,D)$ be a $q$-summable compact Banach spectral triple for some $q>0$. If $S^{q}_{\mathrm{app}}(X)$ has the strong convex compactness property then the proof of Proposition 7.7 shows that the Banach Fredholm module $(X,\pi,\operatorname{\mathrm{sgn}}D)$ is $q$-summable if we can extend the Bochner integral to quasi-Banach spaces.
@@ 8 Examples of triples and Banach Fredholm modules
Suppose that $1<p<\infty$. Here $\mathbb{T}\overset{\mathrm{def}}{=}\{z\in\mathbb{C}:|z|=1\}$ is the one-dimensional torus. The functions defined on $\mathbb{T}$ can be identified with periodic functions on $\mathbb{R}$ with period $2\pi$. In particular, the algebra of continuous complex functions on $\mathbb{T}$ can be written as
$\mathrm{C}(\mathbb{T})=\{f\in\mathrm{C}(\mathbb{R}):f(\theta+2\pi)=f(\theta)\}.$
Recall that by [EnN00, Proposition 4.10 p. 31] any multiplication operator $M_{f}\colon\mathrm{L}^{p}(\mathbb{T})\to\mathrm{L}^{p}(\mathbb{T})$, $g\mapsto fg$ by a measurable function $f\colon\mathbb{T}\to\mathbb{C}$ is bounded if and only if $f$ is essentially bounded. In this case, we have $\left\|M_{f}\right\|_{\mathrm{L}^{p}(\mathbb{T})\to\mathrm{L}^{p}(\mathbb{T})}=\left\|f\right\|_{\mathrm{L}^{\infty}(\mathbb{T})}$. Consequently, we can consider the isometric homomorphism $\pi\colon\mathrm{L}^{\infty}(\mathbb{T})\to\mathrm{B}(\mathrm{L}^{p}(\mathbb{T}))$, $f\mapsto M_{f}$. It is well known, essentially from [Kha13, Exercise 4.2.6 p. 194] and [GVF01, p. 390], that $(\mathrm{C}(\mathbb{T}),\mathrm{L}^{2}(\mathbb{T}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta})$ is a compact spectral triple, where the used homomorphism is the restriction of $\pi$ on the algebra $\mathrm{C}(\mathbb{T})$. First, we prove an $\mathrm{L}^{p}$-generalization of this classical fact. Here, we consider the closure of the closable5554. Let $(f_{n})_{n\geqslant 1}$ be a sequence of functions belonging to $\mathrm{C}^{\infty}(\mathbb{T})$ such that $f_{n}\to 0$ in $\mathrm{L}^{p}(\mathbb{T})$ and $f_{n}^{\prime}\to g$ in $\mathrm{L}^{p}(\mathbb{T})$ for some $g\in\mathrm{L}^{p}(\mathbb{T})$. For any $h\in\mathrm{C}^{\infty}(\mathbb{T})$, integration by parts gives $\int_{0}^{2\pi}f_{n}^{\prime}(\theta)h(\theta)\mathop{}\mathopen{}\mathrm{d}\theta=-\int_{0}^{2\pi}f_{n}(\theta)h^{\prime}(\theta)\mathop{}\mathopen{}\mathrm{d}\theta$. Passing to the limit (using Hölder’s inequality), we obtain $\int_{0}^{2\pi}g(\theta)h(\theta)\mathop{}\mathopen{}\mathrm{d}\theta=0$ for all $h\in\mathrm{C}^{\infty}(\mathbb{T})$. Hence $g=0$ in $\mathrm{L}^{p}(\mathbb{T})$ by [PRR19, Proposition 1.1.5 p. 4]. unbounded operator $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}\colon\mathrm{C}^{\infty}(\mathbb{T})\subset\mathrm{L}^{p}(\mathbb{T})\to\mathrm{L}^{p}(\mathbb{T})$ and we denote again $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ its closure whose the domain $\operatorname{dom}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ is the Sobolev space
@@
Finally, note that the operator $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ is an unbounded Fourier multiplier with symbol $(\mathrm{i}\,n)_{n\in\mathbb{Z}}$.
@@
(8.1) $\left\|f\right\|_{\mathrm{L}^{q}(\mathbb{T}^{d})}\lesssim\left\|f\right\|_{\mathrm{W}^{\alpha,p}(\mathbb{T}^{d})}.$